2. Solve the initial value problem and write y explicitly as a function of a. 4e2 y(0) = 1 y h
Answer #1

2) Given y'=4e^2x/y and y(0)=1

We know that y'=dy/fx

Therefore dy/dx=4e^2x/y

ydy=4e^2xdx

Integrating on both sides , we get

y^2/2 = 2e^2x + c

y^2 = 4e^2x +2c

given y(0)=1 means at x=0 y=1

1=4+2c.

Therefore c=-3/2

So y^2=4e^2x-3/2